OpenStudy (anonymous):

simplify rood of 0.25 X^6

OpenStudy (anonymous):

\[\sqrt{.25x^6} = \sqrt{.25}*\sqrt{x^6}=x^3/2\]

OpenStudy (anonymous):

daniel you are an answering champion. hope it is clear that .25=\[\frac{1}{4}\] and that the square root of 1/4 is 1/2